OpenStudy (anonymous):

solve the equation x bone as it is x^3+5x^2+4 plis

OpenStudy (dumbcow):

\[x ^{3}+5x ^{2}+4 = 0\] ??

OpenStudy (anonymous):

x bone ... right

OpenStudy (dumbcow):

x = -5.15

OpenStudy (anonymous):

x^3+5x^2+4 =x^3+(4+1)x^2+4 =x^3+4x^2+x^2+4 =x^2(x+4)+1(x^2+4) =(x^2)(x+4)(x^2+4)